OpenStudy (anonymous):

what is the derivative of f(x)=x^3-1 at X=3

OpenStudy (anonymous):

x^3-1 3^3 = 9 - 1 = 10

OpenStudy (loser66):

take derivative first, then plug x = 3 into it.